# Enhancing Gravitational Wave Parameter Estimation with Non-Linear Memory: Breaking the Distance-Inclination Degeneracy

### v1

@@ Abstract
In this study, we investigate the role of the non-linear memory effect in gravitational wave (GW) parameter estimation, particularly we explore its capability to break the degeneracy between luminosity distance and inclination angle in binary coalescence events. Motivated by the rapid growth in GW detections and the increasing sensitivity of GW observatories enhancing the precision of cosmological and astrophysical measurements is crucial. We propose leveraging the non-linear memory effect — a subtle, persistent feature in the GW signal resulting from the cumulative impact of emitted gravitational waves — as a novel approach to enhance parameter estimation accuracy. Through a comprehensive series of injection studies, encompassing both reduced and full parameter spaces, we evaluate the effectiveness of non-linear memory in various scenarios for aligned-spin systems. Our findings demonstrate the significant potential of non-linear memory in resolving the inclination-distance degeneracy, particularly for events with high signal-to-noise ratios (SNR $>$ 60) for the current generation of detectors and in the context of future detector sensitivities such as the planned LIGO ${}^{\sharp}$ upgrade. The results also suggest that excluding non-linear memory from parameter estimation could introduce significant systematics in future LIGO ${}^{\sharp}$ detections. This observation will hold even greater weight for next-generation detectors, highlighting the importance of including non-linear memory in GW models for achieving high-accuracy measurements for gravitational wave (GW) astronomy.
@@ I Introduction
However, in many of the LIGO-Virgo compact binary coalescence (CBC) events, the binaries do not have a high mass ratio, precession, or matter effect [1]. As a result, we propose a complementary approach to tackle this issue for binary systems that are symmetric in mass and non-precessing, using the non-linear memory effect. The non-linear memory effect is a permanent spacetime displacement phenomenon that can be interpreted as a gravitational wave component generated by the emission of gravitational waves itself [16]. This effect is not an isolated phenomenon but a crucial element within the broader framework of gravitational wave physics. It stands as one of the pivotal aspects of the ’infrared triangle,’ a conceptual model that intricately links gravitational-wave memory with BMS supertranslations and soft theorems, as explored in depth by Strominger [17]. This interconnection underscores the fundamental nature of gravitational wave memory in understanding the underlying principles of gravitational wave emissions and their implications for theoretical physics. In this work, we limit ourselves to the so-called displacement memory as it is the dominant memory effect which is available for the gravitational waves detectors. Other sub-dominant memory effects such as spin memory[18, 19] and center-of-mass memory [20] are left for future investigations.
@@
The morphology of the non-linear memory is influenced by various factors such as the total mass, mass ratio, and inclination. Predominantly residing in the (2,0) mode for non-precessing systems or the corresponding mode in the co-precessing frame for precessing binaries, the memory effect varies based on the orientation of the system – maximized in edge-on and minimized for face-on systems [23, 24]. This characteristic could serve as a potent tool for disentangling the inclination-distance degeneracy.
Despite its potential, the non-linear memory effect is typically weak and has not been found in the past LIGO-Virgo runs [25, 26]. The prospects of detection of non-linear memory from multiple events with LIGO-Virgo-KAGRA are discussed here [27]. There are a few works to set the detection limit for memory in Einstein Telescope and space-based gravitational wave detectors [28, 29, 30, 31]. Here we aim to investigate how non-linear memory will affect the parameter estimation results through a campaign of injection studies, and understand the parameter range where the inclusion of the memory effect in parameter estimation is necessary for ground-based detectors. We also aspire to understand if failing to consider the memory effect in parameter estimation could lead to systematics. In this work, we have studied the binary systems with aligned spins as generic spins can lead to complex features making it more challenging to study non-linear memory in isolation. Moreover, the non-linear memory peaks at equal mass systems where the effect of spin-precession is expected to be minimal.
This paper is organized as follows. In Section II, we provide a brief explanation of the non-linear memory effect, including its properties. In Section III, we briefly introduce the methods for parameter estimation, the waveform model, and the statistics used in this work. Finally, in Section IV, we present our results obtained from the injection studies including the injections with reduced priors and full priors, as well as using the future ${}^{\sharp}$ upgraded sensitivity, and a selected real event detected by the LVK collaboration.
@@ II Nonlinear memory
The non-linear memory component of gravitational waves can be calculated from the oscillatory modes $h^{lm}$ of gravitational waves for each spherical harmonic ${}_{-2}Y^{lm}$ using the equation [32]
@@
where $R$ denotes the luminosity distance, $c$ represents the speed of light, and $T_{R}$ refers to the retarded time. The term $\dot{h}^{\ell m}$ is the time derivative of the strain, with the overbar indicating its complex conjugate. The coefficient $G_{mm^{\prime}m^{\prime\prime}}^{\ell\ell^{\prime\prime}}$ is defined as follows:
@@
$=(-1)^{m+m^{\prime}}\sqrt{\frac{(2\ell+1)\left(2\ell^{\prime}+1\right)\left(2\ell^{\prime\prime}+1\right)}{4\pi}}$
@@
The above matrices represent the Wigner 3-j symbols, and ${}_{-2}\bar{Y}^{\ell m}$ is the complex conjugate of ${}_{-2}Y^{\ell m}$.
In equation (1), we can see that the memory is a result of combinations of all $(l,m)$ modes. However, it should be noted that each combination contributes with a different amplitude due to the differences in the amplitude and phase of each mode, as well as the varying factor $G_{mm^{\prime}m^{\prime\prime}}^{\ell\ell^{\prime\prime}}$ in equation (2). The amplitude contributions of each combination are displayed in Fig. 2. It is evident that the $(2,2)\times(2,2)$ combination is from 8 times to several orders of magnitude higher than the other modes. Considering the non-linear memory is already around two orders lower than the dominant $(2,2)$ mode, we can safely ignore the contribution from other modes to reduce the computational time significantly.
Non-linear memory contributions alone exhibit characteristics similar to step functions. They increase monotonically and slowly with the compact binary getting closer, then sharply rise at merger, and finally induce a permanent amplitude shift. After a merger event, memory appears as a flat line, indicating its nearly zero frequency as shown in the orange line of the inset in Fig. 1. Ground-based detectors, which are sensitive to tens to a few thousand Hz, cannot detect this “0 Hz” permanent displacement directly, necessitating the application of a high-pass filter to the waveform as shown in the deep blue line in the inset of Fig. 1. The plus polarization waveform contains only (2,$\pm$2) oscillatory modes with/without non-linear memory in $(2,0)$ are shown in Fig. 1 with red and green respectively.
Figure 1: The waveforms shown in the main plot are the plus polarization of gravitational waves (2,$\pm$2) oscillatory modes with and without their corresponding non-linear memory contribution to (2,0) modes. The inset shows the non-linear memory separately in orange and the same signal after a high pass at 10 Hz in dark blue, showing its nearly oscillatory behavior.
Figure 2: The amplitude of memory contribution for different $(l,m)$ mode combinations to $(2,0)$ mode. The blue points show the non-linear memory with positive amplitude while the red points show the ones with negative amplitude. The contribution from $(2,2)$ mode is shown in the star symbol.
Figure 3: The SNR relative fraction of the real part of non-linear memory contribution in (2,0) mode and dominant (2,2) mode in different aligned spin $s_{1z}$ and mass ratio $q$. Here we have used the zero-detuned high-power LIGO noise power spectral density (PSD) as a representative of the current generation of ground-based detectors.
@@ III.1 Bayes parameter estimation
In order to calculate memory accurately from our waveform model, we need to integrate the oscillatory modes precisely. This requires a high sample rate to reduce the error in the integration. However, a higher sample rate can significantly slow down waveform generation. Therefore, we use a different sample rate for each different total mass.
@@ III.3 Relative fractional improvement
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}\left(D_{L}^{i}-\hat{D}_{L}\right).$ (7)
@@
To quantify the detectability of memory, we proposed a new quantity memory SNR $\rho_{20}$ inspired by [40], we can define it as
@@
while the sensitivity of the detector to the $(l,m)$ harmonic, $\sigma_{lm}$, is given by
$\sigma_{lm}=\sqrt{(\tilde{h}_{lm}|\tilde{h}_{lm})}.$ (11)
@@ IV Results
We present a series of injection studies to investigate the effectiveness of non-linear memory in breaking the degeneracy of distance and inclination with a reduced prior, using LIGO-Virgo O5 sensitivity [41]. We then vary all the priors for the non-precessing system to explore the improvement in a more realistic situation. Finally, we discuss the potential improvement that non-linear memory can bring to the distance constraint for the future LIGO ${}^{\sharp}$.
@@ IV.1 Degeneracy breaking for reduced prior
In Fig. 5, we can see the posteriors of two different injection distances, 300 Mpc and 1200 Mpc, in violin plots. These plots demonstrate the enhancement of distance measurement using non-linear memory for high (around 100 at 300 Mpc) and low SNR (around 20 at 1200 Mpc). The upper panel of the plots represents the high SNR case, injected at 300 Mpc, where we can observe that the luminosity distance constraint is much better for face-on and edge-on systems. Additionally, for all inclination angles, the constraint peaks at the injected distance. On the other hand, for the low SNR case, injected at 1200 Mpc, the improvement is not as obvious, but the quartiles represented by the dashed line in the violin plots still show that the recoveries with a non-linear memory model provide better recovery of the injected value.
To quantify the relative fractional improvement of the luminosity distance measurement, we plot the $\Delta D_{L}$ and $\Delta\mathrm{MSE}$ in Fig. 6a and 6b. From Fig. 6a, we can see that the highest improvement is from the closer distance and the face-on system, rather than the edge-on system, where the non-linear memory amplitude is highest, as we might expect. This can be explained by the SNR contour plot overlaid on the $\Delta D_{L}$ contour, which shows the improvement on the $\Delta D_{L}$ following the SNR increase. The relative improvement on the width of the 90% credible interval only provides information about the constraint, but it can’t show the accuracy of the posterior. Therefore, we have also calculated the MSE and plotted the relative fractional improvement of MSE defined in Eq. (6) in Fig. 6b. We observe that even though the SNR at the same distance for the edge-on system is lower than that of the face-on system, due to the maximization of non-linear memory, the $\Delta\mathrm{MSE}$ is good for both systems.
@@ IV.2 Full parameter recovery
We injected the same 25 sets of parameters with varied distances and inclinations as the previously reduced prior section. The contour plot of $\Delta\mathrm{MSE}$ is shown in Fig. 7a. Compared with the same contour plot in Fig. 6b, we can notice that with more parameters open for sampling in the prior, the improvement of the mean-squared error drops for the low SNR scenarios. For the high SNR cases, taken to be, when SNR is greater than 60, the $\Delta\mathrm{MSE}$ shows a similar amount of improvement. We can also notice that in the contour plot near inclination $\pi/4$ and distance $D_{L}$ around 1200 Mpc give worse MSE than the posteriors from the recovery without memory below SNR 30. This can be attributed to the very low (but not negligible) SNR available in the injection for the (2,0) mode which makes the posterior broader when sampling with memory. Also, the contribution from the weak oscillatory part of the (2,0) which is not available in the sampling model, can lead to further worsening of the posterior. This feature is resolved for the edge-on system (inclination of $\pi/2$) as the (2,0) mode will have more SNR.
@@
These results give us hints that for low SNR (SNR $<$ 60), adopting non-linear memory in the parameter estimation won’t gain too much improvement but will have a higher cost on the computing resources with adding additional mode. For the high SNR events, ignoring the non-linear memory effect will introduce significant systematics in the distance-inclination measurement. The $\rho_{20}^{\perp}$ of the injections is greater than 2 for the ones with a luminosity distance greater than 460 Mpc. It is corresponding to SNR 60 when the inclination is $\pi/4$. For current detectors, it is not likely to have such a high SNR and close-distance event. Thus in the next section, we will continue the exploration with the proposed LIGO ${}^{\sharp}$ sensitivity curve [42].
IV.3 Case of LIGO ${}^{\sharp}$

### v2

@@ Abstract
In this study, we investigate the role of the non-linear memory effect in gravitational wave (GW) parameter estimation, particularly we explore its capability to break the degeneracy between luminosity distance and inclination angle in binary coalescence events. Motivated by the rapid growth in GW detections and the increasing sensitivity of GW observatories enhancing the precision of cosmological and astrophysical measurements is crucial. We propose leveraging the non-linear memory effect — a subtle, persistent feature in the GW signal resulting from the cumulative impact of emitted gravitational waves — as a novel approach to enhance parameter estimation accuracy. Through a comprehensive series of injection studies, encompassing both reduced and full parameter spaces, we evaluate the effectiveness of non-linear memory in various scenarios for aligned-spin systems. Our findings demonstrate the significant potential of non-linear memory in resolving the inclination-distance degeneracy, particularly for events with high signal-to-noise ratios (SNR $>$ 90) for the current generation of detectors or closer than 1 Gpc in the context of future detector sensitivities such as the planned LIGO A upgrade. The results also suggest that excluding non-linear memory from parameter estimation could introduce significant systematics in future LIGO A detections. This observation will hold even greater weight for next-generation detectors, highlighting the importance of including non-linear memory in GW models for achieving high-accuracy measurements for gravitational wave (GW) astronomy.
@@ I Introduction
However, in many of the LIGO-Virgo compact binary coalescence (CBC) events, the binaries do not have a high mass ratio, precession, or matter effect [1]. As a result, we propose a complementary approach to tackle this issue for binary systems that are symmetric in mass and non-precessing, using the non-linear memory effect. The non-linear memory (also referred to as null memory [16]) effect is a permanent spacetime displacement phenomenon that can be interpreted as a gravitational wave component generated by the emission of gravitational waves itself [17, 18, 19, 20, 21]. This effect is not an isolated phenomenon but a crucial element within the broader framework of gravitational wave physics. It stands as one of the pivotal aspects of the ’infrared triangle,’ a conceptual model that intricately links gravitational-wave memory with BMS supertranslations and soft theorems, as explored in depth by Strominger [22, 23, 24]. This interconnection underscores the fundamental nature of gravitational wave memory in understanding the underlying principles of gravitational wave emissions and their implications for theoretical physics. In this work, we limit ourselves to the so-called displacement memory as it is the dominant memory effect which is available for the gravitational waves detectors. Other sub-dominant memory effects such as spin memory[25, 26] and center-of-mass memory [27] are left for future investigations.
@@
The morphology of the non-linear memory is influenced by various factors such as the total mass, mass ratio, and inclination. Predominantly residing in the (2,0) mode for non-precessing systems or the corresponding mode in the co-precessing frame for precessing binaries, the memory effect varies based on the orientation of the system – maximized in edge-on and minimized for face-on systems [30]. This characteristic could serve as a potent tool for disentangling the inclination-distance degeneracy.
Despite its potential, the non-linear memory effect is typically weak and has not been found in the past LIGO-Virgo runs [31, 32]. The prospects of detection of non-linear memory from multiple events with LIGO-Virgo-KAGRA are discussed here [33, 34, 35]. There are a few works to set the detection limit for memory in Einstein Telescope and space-based gravitational wave detectors [36, 37, 38]. Additionally, [39] has demonstrated the potential of non-linear memory in resolving the inclination-distance degeneracy for the LISA with Fisher analysis. Here we aim to investigate how non-linear memory will affect the parameter estimation results through a campaign of injection studies, and understand the parameter range where the inclusion of the memory effect in parameter estimation is necessary for ground-based detectors. We also aspire to understand if failing to consider the memory effect in parameter estimation could lead to systematics. In this work, we have studied the binary systems with aligned spins as generic spins can lead to complex features making it more challenging to study non-linear memory in isolation. Moreover, the non-linear memory peaks at equal mass systems where the effect of spin-precession is expected to be minimal.
This paper is organized as follows. In Section II, we provide a brief explanation of the non-linear memory effect, including its properties. In Section III, we briefly introduce the methods for parameter estimation, the waveform model, and the statistics used in this work. Finally, in Section IV, we present our results obtained from the injection studies including the injections with reduced priors and full priors, as well as using the future LIGO-A upgraded sensitivity, and a selected real event detected by the LVK collaboration.
@@ II Nonlinear memory
The non-linear memory component of gravitational waves can be calculated from the oscillatory modes $h^{\ell m}$ of gravitational waves for each spherical harmonic ${}_{-2}Y^{\ell m}$ using the equation [21, 30, 40, 41]
@@
where $R$ denotes the luminosity distance, $c$ represents the speed of light, and $T_{R}$ refers to the retarded time. The term $\dot{h}^{\ell m}$ is the time derivative of the strain, with the overbar indicating its complex conjugate. The coefficient $G_{mm^{\prime}m^{\prime\prime}}^{\ell\ell^{\prime}\ell^{\prime\prime}}$ is defined as follows:
@@
$=(-1)^{m+m^{\prime\prime}}\sqrt{\frac{(2\ell+1)\left(2\ell^{\prime}+1\right)\left(2\ell^{\prime\prime}+1\right)}{4\pi}}$
@@
It is an integral over three spin-weighted spherical harmonics, two with spin weight -2 and one ordinary spherical harmonic function with spin weight 0, which can be expressed in terms of Wigner 3-j symbols. The ${}_{-2}\bar{Y}^{\ell m}$ denote the complex conjugate of ${}_{-2}Y^{\ell m}$.
In equation (1), we can see that the memory is a result of combinations of all $(\ell,m)$ modes. However, it should be noted that each combination contributes with a different amplitude due to the differences in the amplitude and phase of each mode, as well as the varying factor $G_{mm^{\prime}m^{\prime\prime}}^{\ell\ell^{\prime\prime}}$ in equation (2). The amplitude contributions of each combination are displayed in Fig. 1. It is evident that the $(2,2)\times(2,2)$ combination is from 8 times to several orders of magnitude higher than the other modes. Considering the non-linear memory is already around two orders lower than the dominant $(2,2)$ mode, we can safely ignore the contribution from other modes to reduce the computational time significantly.
Non-linear memory contributions alone exhibit characteristics similar to step functions. They increase monotonically and slowly with the compact binary getting closer, then sharply rise at merger, and finally induce a permanent amplitude shift. After a merger event, memory appears as a flat line, indicating its nearly zero frequency as shown in the orange line of the inset in Fig. 2. Ground-based detectors, which are sensitive to tens to a few thousand Hz, cannot detect this “0 Hz” permanent displacement directly. To truly observe the memory effect, one would need freely-falling observers, indicating a permanent, net displacement. This necessitates the application of a high-pass filter to the waveform as shown in the deep blue line in the inset of Fig. 2. The plus polarization waveform contains only (2,$\pm$2) oscillatory modes with/without non-linear memory in $(2,0)$ are shown in Fig. 2 with red and green respectively.
Figure 1: The amplitude of memory contribution for different $(\ell,m)$ mode combinations to $(2,0)$ mode. The blue points show the non-linear memory with positive amplitude while the red points show the ones with negative amplitude. The contribution from $(2,2)$ mode is shown in the star symbol.
Figure 2: The waveforms shown in the main plot are the plus polarization of gravitational waves (2,$\pm$2) oscillatory modes with and without their corresponding non-linear memory contribution to (2,0) modes. The inset shows the non-linear memory separately in orange and the same signal after a high pass at 10 Hz in dark blue, showing its nearly oscillatory behavior. These waveforms are generated using the waveform model described in Sec. III.2 for a total mass 40 $M_{\odot}$, mass ratio 1, non-spinning, and edge-on binary black hole system at 100 Mpc.
Figure 3: The SNR relative fraction of the real part of non-linear memory contribution in (2,0) mode and dominant (2,2) mode in different aligned spin by varying the primary black hole spin in the z-direction, $s_{1z}$, and mass ratio $q$. Here we have used the zero-detuned high-power LIGO noise power spectral density (PSD) as a representative of the current generation of ground-based detectors.
@@ III.1 Bayes parameter estimation
In order to calculate memory accurately from our waveform model, we need to integrate the oscillatory modes precisely. This requires a high sample rate to reduce the error in the integration. However, a higher sample rate can significantly slow down waveform generation. Therefore, we use a different sample rate for each different total mass as shown in Table 1.
Table 1: Sample rates for different total mass range.
Total mass Sample rate
$>80M_{\odot}$ 1024 Hz
$(55M_{\odot},80M_{\odot})$ 2048 Hz
$(25M_{\odot},55M_{\odot})$ 4096 Hz
$(15M_{\odot},25M_{\odot})$ 8192 Hz
$<15M_{\odot}$ 16384 Hz
@@ III.2 Aligned spin waveform model with memory
Figure 5: The mismatch plot between IMRPhenomT and NRHybSur3dq8_CCE (shown in blue), IMRPhenomT_meme and NRHybSur3dq8_CCE (shown in orange). The peaks for each mismatch distribution are indicated by dashed lines.
@@
The figure displayed in Fig. 5 illustrates the mismatch between the waveforms generated by IMRPhenomT and IMRPhenomT_meme as compared to those generated by NRHybSur3dq8_CCE. The mismatch for two waveform $h_{1}$ and $h_{2}$ is defined as
$\textrm{mismatch}(h_{1},h_{2})=1-\max_{\phi,t,\psi}O(h_{1},h_{2})$ (5)
which maximizing the overlap $O(h_{1},h_{2})$ over polarization angles $\psi$, phase $\phi$, and time shifts $t$. The overlap $O(h_{1},h_{2})$ is given by
$O(h_{1},h_{2})=\frac{(h_{1}\mid h_{2})}{\sqrt{(h_{1}\mid h_{1})(h_{2}\mid h_{2})}}$ (6)
and the power spectrum $S(f)$ weighted inner product $(a\mid b)$ is defined as
$(a\mid b):=4\operatorname{Re}\int_{0}^{f_{\max}}\frac{\tilde{a}(f)\tilde{b}(f)^{\star}}{S(f)}\mathrm{d}f$ (7)
We generate the waveforms for each waveform model only with $(2,2)$ and $(2,0)$ modes, and compute the mismatches for binary black holes with total mass 20-80 $M_{\odot}$, mass ratio from 1 to 3, and $s_{1z}$ from 0 to 0.8. The histogram peak indicates a factor of 2 improvement in mismatches for IMRPhenomT_meme vs IMRPhenomT with NRHybSur3dq8_CCE. From this mismatch study, we can see that both IMRPhenomT_meme and the underlying base model IMRPhenomT are highly accurate and sufficient for use in this study. While the relatively small improvement due to the inclusion of the memory effect means it may be difficult to conclusively claim detection of the memory effect at current detector sensitivities, the complementary phenomenology of the memory effect relative to the oscillatory signal makes it perfect for assisting in breaking the distance-inclination degeneracy.
@@ III.3 Relative fractional improvement
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}\left(D_{L}^{i}-\hat{D}_{L}\right)^{2}.$ (10)
@@
Non-linear memory can be considered as one of the higher-order mode multipoles. Thus we can apply the conventional higher-order mode multipole SNR to set out theoretical expectations. Inspired by [49], we proposed a new quantity memory SNR $\rho_{20}$ , we can define it as
@@
while the sensitivity of the detector to the $(\ell,m)$ harmonic, $\sigma_{\ell m}$, is given by
$\sigma_{\ell m}=\sqrt{(\tilde{h}_{\ell m}|\tilde{h}_{\ell m})}.$ (14)
where $\tilde{h}_{\ell m}$ is the higher order harmonics $(\ell,m)$ waveform in the frequency domain
@@ IV Results
We present a series of injection studies to investigate the effectiveness of non-linear memory in breaking the degeneracy of distance and inclination with a reduced prior, using LIGO-Virgo O5 sensitivity [50]. We then vary all the priors for the non-precessing system to explore the improvement in a more realistic situation. Finally, we discuss the potential improvement that non-linear memory can bring to the distance constraint for the future LIGO A.
@@ IV.1 Degeneracy breaking for reduced prior
In Fig. 6, we can see the posteriors of two different injection distances, 300 Mpc and 1200 Mpc, in violin plots. These plots demonstrate the enhancement of distance measurement using non-linear memory for high (around 160 at 300 Mpc) and low SNR (around 40 at 1200 Mpc). The upper panel of the plots represents the high SNR case, injected at 300 Mpc, where we can observe that the luminosity distance constraint is much better for face-on and edge-on systems. Additionally, for all inclination angles, the constraint peaks at the injected distance. On the other hand, for the low SNR case, injected at 1200 Mpc, the improvement is not as obvious, but the quartiles represented by the dashed line in the violin plots still show that the recoveries with a non-linear memory model provide better recovery of the injected value.
To quantify the relative fractional improvement of the luminosity distance measurement, we plot the $\Delta D_{L}$ and $\Delta\mathrm{MSE}$ in Fig. 7a and 7b. From Fig. 7a, we can see that the highest improvement is from the closer distance and the face-on system, rather than the edge-on system, where the non-linear memory amplitude is highest, as we might expect. This can be explained by the SNR contour plot overlaid on the $\Delta D_{L}$ contour, which shows the improvement on the $\Delta D_{L}$ following the SNR increase. The relative improvement on the width of the 90% credible interval only provides information about the constraint, but it can’t show the accuracy of the posterior. Therefore, we have also calculated the MSE and plotted the relative fractional improvement of MSE defined in Eq. (9) in Fig. 7b. We observe that even though the SNR at the same distance for the edge-on system is lower than that of the face-on system, due to the maximization of non-linear memory, the $\Delta\mathrm{MSE}$ is good for both systems.
@@
We further explored injections with similar parameters, varying the aligned spin magnitude $s_{1z}$, as detailed in Appendix A. The findings indicate that when the spin magnitude $s_{1z}$ is less than 0.6, the enhancements attributable to memory effects remain largely consistent. Consequently, our focus on cases with zero spin is representative of most scenarios involving aligned spin.
@@ IV.2 Full parameter recovery
We injected the same 25 sets of parameters with varied distances and inclinations as the previously reduced prior section. The contour plot of $\Delta\mathrm{MSE}$ is shown in Fig. 8a. Compared with the same contour plot in Fig. 7b, we can notice that with more parameters open for sampling in the prior, the improvement of the mean-squared error drops for the low SNR scenarios. For the high SNR cases, taken to be, when SNR is greater than 90, the $\Delta\mathrm{MSE}$ shows a similar amount of improvement. We can also notice that in the contour plot near inclination $\pi/3$ and distance $D_{L}$ around 1200 Mpc give worse MSE than the posteriors from the recovery without memory below SNR 40. This can be attributed to the very low (but not negligible) SNR available in the injection for the (2,0) mode which makes the posterior broader when sampling with memory. Also, the contribution from the weak oscillatory part of the (2,0) which is not available in the sampling model, can lead to further worsening of the posterior. This feature is resolved for the edge-on system (inclination of $\pi/2$) as the (2,0) mode will have more SNR.
@@
These results give us hints that for low SNR (SNR $<$ 90), adopting non-linear memory in the parameter estimation won’t gain too much improvement but will have a higher cost on the computing resources with adding additional mode for LIGO O5. For the high SNR events, ignoring the non-linear memory effect will introduce significant systematics in the distance-inclination measurement. The $\rho_{20}^{\perp}$ of the injections is greater than 2.3 for the ones with a luminosity distance closer than 550 Mpc. It is corresponding to SNR 90 when the inclination is $\pi/4$. For current detectors, it is not likely to have such a high SNR and close-distance event. Thus in the next section, we will continue the exploration with the proposed LIGO A sensitivity curve [51].
IV.3 Case of LIGO A